OpenStudy (anonymous):

Please verify, find the x intercept of the function f(x)=2 ln(x-3) 0=2(lnx-3) 0=lnx-3 3=lnx 3= loge^x X= e^'3

OpenStudy (mathmale):

0=2(lnx-3) OK 0=lnx-3 Should keep those parentheses! 0=ln(x-3) 3=lnx 3= loge^x There is no advantage to re-writing 3=ln x. 3=ln x is correct as is. X= e^'3 Mind explaining how you got this result? Please delete the ( ' )

OpenStudy (anonymous):

I replaced x with 3 to get e^3 or should it be -3

OpenStudy (mathmale):

(In answer to your question: No.) If 3=lnx: Treat both sides of this equation as exponents. Write the simple equation e = e Use that 3 as the exponent for the first e and that ln x as the exponent for the second e. Write this out, please.

OpenStudy (anonymous):

e^3 = e^3? I'm not sure what you want me to do. Or is the answer ln^ 3 = loge^3 X inltercept loge^3

Parth (parthkohli):

\[\ln(x - 3) =0\]\[\log_{e}{(x-3)} =0\]\[e^0 = x - 3\]

Parth (parthkohli):

Note that \(\ln(x - 3)\) should NOT be written as \(\ln(x) - 3\).

Parth (parthkohli):

Oh, just noticed that mathmale had already pointed that out. Sorry.

OpenStudy (mathmale):

@parthkohli: Here's your medal for noticing what I had not noticed! @Nancyo: Sorry. Looks like it's back to the drawing board. We'll need to start over. Solve for x, if 0=2ln(x-3).

OpenStudy (mathmale):

Think: ln 1 = 0. Can you use this fact to solve 0=2ln(x-3)?

OpenStudy (anonymous):

I may be way off base Ln 1 =0 bc e^0 =1 Ln e =1 bc e^1 =e So now I have: 0= 2ln(lnx-3) 0= ln(x-3) Loge(x-3)=0 e= x-3. 1=x-3 X=4?

OpenStudy (mathmale):

ln 1 = 0. Can you use this fact to solve 0=2ln(x-3)? Here's how: We know that ln 1 = 0. Therefore, if ln(x-3) = 0, x-3 must be 1. So your x=4 is correct. any questions regarding this particular math problem?

OpenStudy (anonymous):

I think I understand now, thank you very much for the help!